# Hyperon spin correlation in high-energy heavy-ion collisions

### v1

@@ Introduction.
The hot and dense matter produced in non-central heavy-ion collisions, known as the quark-gluon plasma (QGP), carries a large amount of orbital angular momentum (OAM) in the form of a fluid velocity field with nonvanishing vorticity. Quarks in the QGP are shown to be globally polarized through the spin-orbital coupling. It was predicted that this leads to a global polarization of the final-state hyperons [1, 2, 3, 4]. Such a phenomenon has been confirmed by the STAR experiment at the Relativistic Heavy-ion Collider (RHIC) [5, 6, 7, 8] as well as by the HADES experiment at lower energies [9], while the global polarization at the LHC energies is close to zero [10] as expected. Recently, an interesting azimuthal and rapidity dependence of the hyperon polarization has also been observed [11, 12], which is related to the local structure of the vorticity or the fluid shear in the QGP [13, 14, 15, 16, 17, 18, 19, 20, 21].
The OAM in heavy-ion collisions was also predicted to lead to the spin alignment of vector mesons [22], which was indeed observed recently for $\phi$ mesons [23]. However, theoretical predictions for the spin alignment induced by conventional mechanisms such as the fluid vorticity or the magnetic field [24, 25, 26, 27] are almost two orders of magnitude smaller than what was observed in the experiments. To account for the disparity, several new mechanisms have been proposed [25, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Among them is the local strong force field (SFF) [25, 30, 31, 35], which can polarize quarks through its coupling to the quark spin. This field is dominated by fluctuations and thus its mean value over a large space and time scale should be zero. As a consequence, it does not contribute to the global polarization of quarks and final-state hyperons. However, since the vector meson’s spin alignment is related to the spin correlation between the constituent quark and anti-quark [22, 24], the fluctuating SFF can cause the spin alignment if its field correlation is nonzero over the size of the vector meson. This model is shown to describe the transverse-momentum dependence of the observed $\phi$ meson spin alignment and predicts the azimuthal-angle or rapidity dependence which can be tested in future experiments [30, 35].
@@
Recently the STAR collaboration has observed a significant $\Lambda\bar{\Lambda}$ spin correlation in proton-proton collisions [41]. This correlation is dominated by the hadronization of entangled $s\bar{s}$ pairs produced via virtual gluon splitting, i.e., $g\rightarrow s+\bar{s}\rightarrow\Lambda+\bar{\Lambda}+X$. Due to helicity conservation, the spin orientations of $s$ and $\bar{s}$ are enforced to align in the same direction. This maximizes the spin correlation of $s\bar{s}$, which is in turn transferred to the final-state $\Lambda\bar{\Lambda}$ spin correlation [41, 42, 43]. By contrast, in nucleus-nucleus collisions, the $\Lambda\bar{\Lambda}$ pairs are from the hadronization of QGP, while their constituent quarks originate from different processes and thus are unlikely to be entangled. The spin correlation arises only if both hyperons are polarized by the same external field in the surrounding hot medium. Such a kind of correlation corresponds to the “induced spin correlation” in Refs. [44, 45].
@@
According to the non-relativistic coalescence model [39, 24, 46], the polarization of a $\Lambda$ baryon is fully carried by its constituent $s$ quark, ${\bf P}_{\Lambda}({\bf x},{\bf p})=\left\langle{\bf P}_{s}({\bf x}_{s},{\bf p}_{s})\right\rangle_{\Lambda}$, where ${\bf P}_{\Lambda}$ and ${\bf P}_{s}$ are the non-relativistic polarizations of a $\Lambda$ hyperon and an $s$ quark, respectively. Here, $\left\langle\cdots\right\rangle_{\Lambda}$ denotes an average over the position ${\bf x}_{s}$ and momentum ${\bf p}_{s}$ of the $s$ quark, weighted by the wave function of the $\Lambda$. For a relativistic $\Lambda$ hyperon, we generalize the above relation to the following form
@@
where $R_{s}\equiv m_{s}/m_{\Lambda}$ is the ratio between the invariant mass of the $s$ quark and the $\Lambda$. Here we have neglected the size of the $\Lambda$ and the motion of the $s$ quark relative to the $\Lambda$ hyperon, indicating that the $\Lambda$ hyperon and the constituent $s$ quark have the same position and velocity. The polarization of $\bar{\Lambda}$ is related to that of $\bar{s}$ in the same way. The relation Eq. (3) has been previously used in Ref. [21] as the “strange-memory scenario” to study the local polarization of $\Lambda$.
@@
where $d\Sigma_{\mu}(x)$ is the differential volume vector at point $x$ on the freezeout hypersurface and the hyperons’ polarization $P^{\mu}_{\Lambda/\bar{\Lambda}}$ is related to the quark polarization [Eq. (4)] through Eq. (3). If we also implement the local-equilibrium condition for $\Lambda$ ($\bar{\Lambda}$) hyperons, the distribution $f_{\Lambda/\bar{\Lambda}}$ should be the Fermi-Dirac distribution
@@
where $\mu_{B}$ is the baryon chemical potential. Analogously, the spin correlation in Eq. (1) can be calculated via
@@
which denotes the difference between the “same-sign” pairs ($\Lambda\Lambda$ or $\bar{\Lambda}\bar{\Lambda}$) and the “opposite-sign” pairs ($\Lambda\bar{\Lambda}$). We show in Figs. 3 and 4 its dependence on the azimuthal-angle difference and the rapidity difference between the two hyperons. When the SFF is taken into account, the $xx$, yyyy, and $zz$ components of $c_{\text{net}}$ peak at $\delta\phi=0$ or $\delta Y=0$ and approach zero when $\delta\phi$ or $\delta Y$ increase, as indicated by the orange, red, and cyan lines. In the absence of SFF, $c_{\text{net}}$ becomes 2-3 orders of magnitude smaller than in the case with SFF, shown in the insets in Figs. 3 and 4. This behavior arises because the SFF has an opposite effect on the polarizations of $\Lambda$ and $\bar{\Lambda}$, as described by Eq. (5), leading to opposite contributions to the spin correlations of $\Lambda\Lambda$ and $\Lambda\bar{\Lambda}$. In contrast, the polarizations induced by the thermal vorticity and the thermal shear tensor are identical for $\Lambda$ and $\bar{\Lambda}$, resulting in nearly the same values of $c_{\Lambda\Lambda}^{yy}$, $c_{\bar{\Lambda}\bar{\Lambda}}^{yy}$, and $c_{\Lambda\bar{\Lambda}}^{yy}$ when the SFF is switched off. The spin Hall effect could also result in a nonzero $c_{\text{net}}$, but its magnitude is much smaller than the contribution of SFF. Therefore $c_{\text{net}}$ is shown to be sensitive to the SFF and is almost independent from the other hydrodynamic fields.
@@
The dependence of the net spin correlation on $\delta\phi$ or $\delta Y$ is a consequence of the position-momentum correlation of hyperons. Due to the transverse expansion of the QGP, the mean transverse position of hyperons emitted from the freeze-out hypersurface is approximately aligned with the transverse momentum, $\left\langle{\bf x}_{T}\right\rangle\propto\hat{\bf p}_{T}$. Similarly, the longitudinal expansion leads to a nearly linear dependence of $\left\langle z\right\rangle$ on the rapidity. Given such significant position-momentum correlations, a larger separation of two hyperons in momentum space corresponds to a larger distance between their production positions, which leads to weaker field-field and spin-spin correlations according to Eqs. (8) and (9). The $\delta\phi$ or $\delta Y$ dependence is similar to the $\Lambda\bar{\Lambda}$ spin correlation observed in p+p collisions [41], but arises from a different mechanism. In p+p collisions, it is dominated by gluon splittings $g\rightarrow\Lambda+\bar{\Lambda}+X$, while the correlation in this study is dominated by interactions between constituent quarks and the fluctuating SFF.

### v2

@@ Introduction.
The hot and dense matter produced in non-central heavy-ion collisions, known as the quark-gluon plasma (QGP), carries a large amount of orbital angular momentum (OAM) in the form of a fluid velocity field with nonvanishing vorticity. Quarks in the QGP are shown to be globally polarized through the spin-orbital coupling. It was predicted that this leads to a global polarization of the final-state hyperons [1, 2, 3, 4]. Such a phenomenon has been confirmed by the STAR experiment at RHIC [5, 6, 7, 8] as well as by the HADES experiment at lower energies [9], while the global polarization at the LHC energies is close to zero [10] as expected. Recently, an interesting azimuthal and rapidity dependence of the hyperon polarization has also been observed [11, 12], which is related to the local structure of the vorticity or the fluid shear in the QGP [13, 14, 15, 16, 17, 18, 19, 20, 21].
The OAM in heavy-ion collisions was also predicted to lead to the spin alignment of vector mesons [22], which was indeed observed recently for $\phi$ mesons [23]. However, theoretical predictions for the spin alignment induced by conventional mechanisms such as the fluid vorticity or the magnetic field [24, 25, 26, 27] are almost two orders of magnitude smaller than what was observed in the experiments. To account for the disparity, several new mechanisms have been proposed [25, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. Among them is the local strong force field (SFF) [25, 30, 31, 35], which can polarize quarks through its coupling to the quark spin. This field is dominated by fluctuations and thus its mean value over a large space and time scale should be zero. As a consequence, it does not contribute to the global polarization of quarks and final-state hyperons. However, since the vector meson’s spin alignment is related to the spin correlation between the constituent quark and anti-quark [22, 24], the fluctuating SFF can cause the spin alignment if its field correlation is nonzero over the size of the vector meson. This model is shown to describe the transverse-momentum dependence of the observed $\phi$ meson spin alignment and predicts the azimuthal-angle or rapidity dependence which can be tested in future experiments [30, 35].
@@
Recently the STAR collaboration has observed a significant $\Lambda\bar{\Lambda}$ spin correlation in proton-proton collisions [43]. This correlation is dominated by the hadronization of entangled $s\bar{s}$ pairs produced via virtual gluon splitting, i.e., $g\rightarrow s+\bar{s}\rightarrow\Lambda+\bar{\Lambda}+X$. Due to helicity conservation, the spin orientations of $s$ and $\bar{s}$ are enforced to align in the same direction. This maximizes the spin correlation of $s\bar{s}$, which is in turn transferred to the final-state $\Lambda\bar{\Lambda}$ spin correlation [43, 44, 45, 27, 46]. By contrast, in nucleus-nucleus collisions, the $\Lambda\bar{\Lambda}$ pairs are from the hadronization of QGP, while their constituent quarks originate from different processes and thus are unlikely to be entangled. The spin correlation arises only if both hyperons are polarized by the same external field in the surrounding hot medium. Such a kind of correlation corresponds to the “induced spin correlation” in Refs. [47, 48].
@@
According to the non-relativistic coalescence model [41, 24, 49], the polarization of a $\Lambda$ baryon is fully carried by its constituent $s$ quark. We generalize such a relation to the relativistic case by setting
@@
where $P_{\Lambda}^{\mu}$ and $P_{s}^{\mu}$ are polarizations for a $\Lambda$ hyperon and an $s$ quark, respectively, and $R_{s}\equiv m_{s}/m_{\Lambda}$ is the ratio between their invariant masses. Here we have neglected the size of the $\Lambda$ and the motion of the $s$ quark relative to the $\Lambda$ hyperon, indicating that the $\Lambda$ hyperon and the constituent $s$ quark have the same position and velocity. The polarization of $\bar{\Lambda}$ is related to that of $\bar{s}$ in the same way. The relation Eq. (3) has been previously used in Ref. [21] as the “strange-memory scenario” to study the local polarization of $\Lambda$.
@@
where $d\Sigma_{\mu}(x)$ is the differential volume vector at point $x$ on the freezeout hypersurface and $P^{\mu}_{\Lambda/\bar{\Lambda}}$ is related to the quark polarization [Eq. (4)] through Eq. (3). If we also implement the local-equilibrium condition for $\Lambda$ ($\bar{\Lambda}$) hyperons, the distribution $f_{\Lambda/\bar{\Lambda}}$ should be the Fermi-Dirac distribution
@@
where $\mu_{B}$ is the baryon chemical potential. Analogously, the spin correlation in Eq. (1) is calculated via
@@
which denotes the difference between the “same-sign” pairs ($\Lambda\Lambda$ or $\bar{\Lambda}\bar{\Lambda}$) and the “opposite-sign” pairs ($\Lambda\bar{\Lambda}$). We show in Figs. 3 and 4 its dependence on the azimuthal-angle difference and the rapidity difference between the two hyperons. When the SFF is taken into account, the $xx$, yyyy, and $zz$ components of $c_{\text{net}}$ peak at $\delta\phi=0$ or $\delta Y=0$ and approach zero when $\delta\phi$ or $\delta Y$ increase. In the absence of SFF, $c_{\text{net}}$ becomes 2-3 orders of magnitude smaller than in the case with SFF, shown in the insets in Figs. 3 and 4. This behavior arises because the SFF has an opposite effect on the polarizations of $\Lambda$ and $\bar{\Lambda}$, as described by Eq. (5), leading to opposite contributions to the spin correlations of $\Lambda\Lambda$ and $\Lambda\bar{\Lambda}$. In contrast, the polarizations induced by the thermal vorticity and the thermal shear tensor are identical for $\Lambda$ and $\bar{\Lambda}$, resulting in nearly the same values of $c_{\Lambda\Lambda}^{yy}$, $c_{\bar{\Lambda}\bar{\Lambda}}^{yy}$, and $c_{\Lambda\bar{\Lambda}}^{yy}$ when the SFF is switched off. The spin Hall effect could also result in a nonzero $c_{\text{net}}$, but its magnitude is much smaller than the contribution of SFF. Therefore $c_{\text{net}}$ is shown to be sensitive to the SFF and is almost independent from the other hydrodynamic fields.
@@
The dependence of the net spin correlation on $\delta\phi$ or $\delta Y$ is a consequence of the position-momentum correlation of hyperons. Due to the transverse expansion of the QGP, the mean transverse position of hyperons emitted from the freeze-out hypersurface is approximately aligned with the transverse momentum, $\left\langle{\bf x}_{T}\right\rangle\propto\hat{\bf p}_{T}$. Similarly, the longitudinal expansion leads to a nearly linear dependence of $\left\langle z\right\rangle$ on the rapidity. Given such significant position-momentum correlations, a larger separation of two hyperons in momentum space corresponds to a larger distance between their production positions, which leads to weaker field-field and spin-spin correlations according to Eqs. (8) and (9). The $\delta\phi$ or $\delta Y$ dependence is similar to the $\Lambda\bar{\Lambda}$ spin correlation observed in p+p collisions [43], but arises from a different mechanism. In p+p collisions (or for high-momentum hyperons in A+A collisions [59, 60]), it is dominated by the parton fragmentation, while the correlation in this study is dominated by interactions between constituent quarks and the fluctuating SFF.